# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v2

@@ Conclusion.
Our initial experiments in line with related works, demonstrate that latent representations computed using the DDIM inversion deviate from the expected characteristics of independent Gaussian noise. Specifically, both visual analysis and quantitative evaluations reveal significant correlations between the neighboring pixels.
@@ 3.3 Location of latent encodings space
To answer it, we first investigate the geometric aspect of how the latent encodings are distributed in relation to the original noise and generated samples. We focus on understanding their spatial relationships across different model architectures and configurations.
First, in Tab. 3, we determine the most probable angles formed by noises ${\mathbf{x_{T}}}$, samples ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ for each model by binning the empirically observed angles and selecting those that satisfy triangle angle sum, while maximizing the joint probability. We provide details on this methodology and visualization in Appendix B. Results of the experiment show that the angle located at the image and noise vertices (accordingly $\angle{\mathbf{x_{0}}}$ and $\angle{\mathbf{x_{T}}}$) are always acute and, in almost every case, the angle by the latent vertex ($\angle{\mathbf{\hat{x}_{T}}}$) is obtuse. This property implies that, due to approximation errors in the reverse DDIM process, latents reside in proximity to, but with a measurable offset, from the shortest-path trajectory between the noise distribution and the generated image.
@@
Table 3: Most probable angles next to the Gaussian noise ${\mathbf{x_{T}}}$, generations ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ across different diffusion models. We use $T=100$ denoising and noising steps.
However, the trajectory defined by the backward diffusion process does not necessarily follow the linear path, as DMs learn a score function that is highly non-linear, causing the denoising path to follow the curved structure of the data manifold [45]. Therefore, to further strengthen our analysis, we investigate the location of the latent with respect to the generation trajectory. To that end, we analyze the distance between the next steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the appropriate latent – along the $\overline{{\mathbf{x_{T}}}{\mathbf{\hat{x}_{T}}}}$ side of the noise-sample-latent triangle. We present the results of this experiment in Fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$.
Figure 3: $l_{2}$-distances between intermediate diffusion denoising steps $x_{t}$ and points on the linear interpolation path from noise ${\mathbf{x_{T}}}$ to the inverted latents ${\mathbf{\hat{x}_{T}}}$. The consecutive intermediate generation steps along the sampling trajectory consequently get closer to the latent variables.
We observe that, while moving from the random Gaussian noise (${\mathbf{x_{T}}}$) towards the final sample (${\mathbf{x_{0}}}$), the intermediate steps $x_{t}$ approach the latent (${\mathbf{\hat{x}_{T}}}$), while after the transition point present around 50-70% of the generative trajectory the distance to the latent becomes lower than the distance to the starting noise. This observation reveals, that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
To more closely localize the latent encodings, we propose to study their relations to the center of the analyzed space - point 00 being the mean of the initial noise ${\mathbf{x_{T}}}$. In Tab. 4, we show the average variance of the latent encodings for randomly sampled and inverted 10,000 images. Latent encodings ${\mathbf{\hat{x}_{T}}}$ have, in general, lower variance than the noise, which means that the inversion process drives them towards the mean of the Gaussian distribution. The remaining question is how does it manifest?
@@
The visualizations shown in Fig. 2 indicate that, in some cases, we can distinguish the structure of the original objects in the latent encodings. In particular, while the pixels associated with the objects have high diversity, which is also expected for white Gaussian noise, the areas related to the background are much smoother.
This leads to the hypothesis that the most significant difference between the initial Gaussian noises and latent encodings is the limited variance in the areas corresponding to the background of the generated image. To validate it, we compare the properties of resulting latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$ that select smooth surfaces in the image, according to the neighbor differences as described in Appendix C. In Tab. 5, we show that the error between the starting noise ${\mathbf{x_{T}}}$ and the latent ${\mathbf{\hat{x}_{T}}}$ resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across all the models, this trend goes along with a decrease in the standard deviation of latents’ pixels related to those image regions. It suggests that DDIM inversion struggles with reversing the denoising operation for those image parts, bringing them to 0 (average) and reducing their diversity.
@@
Table 5: Average per-pixel absolute error (between noise and latent encodings) and standard deviation of the latents’ pixels corresponding to the plain and non-plain areas in images. The error for plain pixels, where latents are less diverse, is higher than for other regions.
@@ Conclusion.
Latent encodings resulting from the DDIM Inversion deviate from the Gaussian noises towards zero values. This can be attributed to the less diverse values for parts of the latents corresponding to the plain image surfaces.
@@ 3.4 Origin of the divergence
Given the observation from the previous section, we now investigate the possible source of the differences between latents and their corresponding noises, posing the question:
Research Question 3: What is the reason behind the specific divergence between input noise and latent calculated with DDIM inversion?
@@
We first recall that the DDIM Inversion error can be attributed to the approximation of the diffusion model’s output at step $t\in{1\dots T}$ with the output from step $t-1$ (see Eq. 4). Hence, we can define the inversion approximation error for step $t$ as the difference between DMs output for the target and previous timesteps ${t}$ and ${t-1}$ as:
where $\mathcal{E}_{t}^{S}$ is the ground truth DM output for inversion step while $\mathcal{E}_{t}^{I}$ is the approximated prediction.
Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the following diffusion steps. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To ascertain that we measure the error solely for the analyzed step $t$, we start the inversion procedure from the ground truth value for step $(t+1)$ assuming perfect inversion for each timestep above $t$. We split the latent pixels into plain and non-plain areas according to the masks calculated for final generations. We explain this setup in detail in Appendix D.
@@
Table 6: Sum of the normalized diffusion model prediction error ¯$\bar{\mathcal{E}}_{(t_{s},t_{e})}^{E}$ during the DDIM inversion for plain and non-plain pixels. Inversion approximation error is higher for the pixels related to the plain image areas, especially in the first 10% inversion steps that contribute to the error more than the remaining 90% of steps. Values in the table are multiplied by $10^{4}$.
In Fig. 4, we present the visualization of calculated differences for each diffusion step. As visible, there is a significant difference in the prediction errors for plain and non-plain areas, especially in the initial steps of the noising process. Additionally, in Tab. 6, we show the actual values of the approximation error. By splitting the inversion process into two parts, we can observe that the first 10% of inversion steps contribute more to the final error than the remaining 90%. Notably, the error in plain areas predominantly accumulates within this first 10% of the inversion steps.
@@ Conclusion.
Differences between noises and latents can be attributed to the higher approximation error for the pixels related to the plain image areas, especially in the first DDIM inversion steps.
4 Consequences of the differences
After identifying the differences between the Gaussian noises and latents and highlighting the origin of the phenomenon, we finally pose the last question:
Research Question 4: What are the practical consequences of the divergence between noises and inverse DDIM latents?
Based on the preceding observations, we hypothesize that the less diverse areas in latent variables, connected to smoothened regions in the images, negatively impact their applicability as priors for the generative process. In this section, we showcase several consequences of this problem, including worse performance in the task of image interpolation and editing, collapsing diversity of generations, and inconsistencies in noise-image mapping.
@@ 4.1 Interpolation quality
We first adhere to the task of image interpolation, where, for two given images, the goal is to generate a sequence of semantically meaningful intermediary frames between them. Classic techniques based on the pixels’ morphing [52] fail to produce content beyond provided inputs [55], leading to outputs with unnatural artifacts. Therefore, numerous methods [7, 44, 41, 58, 55, 1] leverage diffusion models with noise representations of images in order to obtain realistic transitions between them. Song et al. [44] propose to use the latent variables resulting from the DDIM inversion technique and interpolate them with the spherical linear interpolation (SLERP [42]), that, for two objects $x$ and $y$, is defined as
$z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y,$ (6)
where $\lambda\in[0;1]$ is an interpolation coefficient and $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
@@
We compare the quality of interpolations in the noise space and the latent space. To this end, we sample $N=20$k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into the latents. Next, we randomly assign pairs of either two noises or latents, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In Fig. 5, we show that, by calculating FID-10k to the training set, intermediate noise generations preserve consistent quality along the entire interpolation path. In contrast, this quality collapses for the setting with latents as inputs, especially when both of them contribute equally to the intermediate interpolation point.
In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall [24], especially in the middle of the interpolation path.
@@
Figure 5: The quality of images generated from spherical linear interpolations of latents ${\mathbf{\hat{x}_{T}}}$ deteriorates along the interpolation path, as indicated by the FID peak in the middle. In contrast, the quality of generations from noise ${\mathbf{x_{T}}}$ interpolations remains stable. Recall metric reveal that the use of latent interpolation leads to the generation of much less diverse than noise.
4.2 Diversity in conditional generation
Text-to-image diffusion models are often combined with deterministic schedulers to manipulate the samples by starting from the same noise and modifying the denoising trajectory with a different prompt. This method, combined with inversion, is widely used to modify real images [12, 32, 10, 17]. However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting state for backward diffusion limits the edited generations.
To evaluate this, we leverage DiT [36] and IF [46] as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected prompts (ImageNet classes for DiT, COCO-30K [26] prompts for IF): (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used during editing.

### v3

@@ Conclusion.
Recent studies, supported by our initial experiments, demonstrate that latent representations computed using the DDIM inversion deviate from the expected characteristics of independent Gaussian noise. Specifically, both visual analysis and quantitative evaluations reveal significant correlations between the neighboring pixels.
@@ 3.2 Location of latent encodings space
To answer this question, we first investigate the geometric aspect of how the latent encodings are distributed in relation to the original noise and generated samples. In table 3, we determine the most probable angles formed by noises ${\mathbf{x_{T}}}$, samples ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ for each model by binning the empirically observed angles and selecting those that satisfy the triangle angle sum, while maximizing the joint probability. We provide details on this methodology and visualizations in appendix F. Results of the experiment show that the angle located at the image and noise vertices (accordingly $\angle{\mathbf{x_{0}}}$ and $\angle{\mathbf{x_{T}}}$) are always acute and, in almost every case, the angle by the latent vertex ($\angle{\mathbf{\hat{x}_{T}}}$) is obtuse. This property implies that, due to approximation errors in the reverse DDIM process, latents reside in proximity to, but with a measurable offset, from the shortest-path trajectory between the noise distribution and the generated image.
@@
Table 3: Most probable angles next to the Gaussian noise ${\mathbf{x_{T}}}$, generations ${\mathbf{x_{0}}}$, and latents ${\mathbf{\hat{x}_{T}}}$ across different diffusion models.
Figure 3: $l_{2}$-distances between intermediate diffusion denoising steps $x_{t}$ and points on the linear interpolation path from noise ${\mathbf{x_{T}}}$ to the inverted latents ${\mathbf{\hat{x}_{T}}}$. The consecutive intermediate generation steps along the sampling trajectory consequently approach the latent.
However, the trajectory defined by the backward diffusion process does not necessarily follow the linear path, as DMs learn a highly non-linear score function, causing the denoising path to follow the curved structure of the data manifold [51]. Therefore, to strengthen our analysis further, we investigate the location of the latent with respect to the generation trajectory. To that end, we analyze the distance between the following steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the associated latent. We present the results of this experiment in fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$.
@@
We observe that, while moving from the random Gaussian noise (${\mathbf{x_{T}}}$) towards the final sample (${\mathbf{x_{0}}}$), the intermediate steps $x_{t}$ approach the latent (${\mathbf{\hat{x}_{T}}}$), while after the transition point present around 50-70% of the generative trajectory the distance to the latent becomes lower than the distance to the starting noise. This observation reveals that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
To more closely localize the latent encodings, we propose to study their relations to the center of the analyzed space, point 00 being the mean of the initial noise ${\mathbf{x_{T}}}$. In table 4, we show the average variance of the latent encodings for randomly sampled and inverted 10,000 images. Latent encodings ${\mathbf{\hat{x}_{T}}}$ have, in general, lower variance than the noise, which means that the inversion process drives them towards the mean of the Gaussian distribution. The remaining question is, how does it manifest?
The visualizations shown in fig. 2 indicate that, in some cases, we can distinguish the structure of the original objects in the latent encodings. In particular, while the pixels associated with the objects have high diversity, which is also expected for white Gaussian noise, the areas related to the background are much smoother. This finding leads to the hypothesis that the most significant difference between the initial Gaussian noises and latent encodings is the limited variance in the areas corresponding to the background of the generated images. To validate it, we compare the properties of the resulting latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$
@@
Table 5: Average per-pixel absolute error (between noise and latent) and standard deviation of the latents’ pixels corresponding to the plain and non-plain areas of generations. The error for plain pixels, where latents are less diverse, is higher than for other regions.
that select smooth surfaces in the image, according to the neighbor differences described in appendix G. In table 5, we show that the error between the starting noise ${\mathbf{x_{T}}}$ and the latent ${\mathbf{\hat{x}_{T}}}$ resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across all the models, this trend goes along with a decrease in the standard deviation of the latents’ pixels related to those image regions. It suggests that DDIM inversion struggles with reversing the plain image areas, bringing them to 0 (average) and reducing their diversity.
@@ Conclusion.
Latent encodings resulting from the DDIM Inversion deviate from the Gaussian noise towards zero values. This can be attributed to the less diverse values for parts of the latents corresponding to the plain image surfaces. This observation reveals that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
@@ 3.3 Origin of the divergence
Given the observation from the previous section, we now investigate the source of the correlations occurring in latent encodings, posing the question:
Research Question 3: What causes the spatial correlations observed in DDIM latents?
@@
Table 6: Total per-pixel inversion error (normalized) over different timestep ranges ¯$\bar{\mathcal{E}}_{(t_{s},t_{e})}^{E}$ for plain and non-plain areas. Inversion approximation error is higher for pixels related to plain image areas, especially in the first 10% of inversion steps.
We first recall that the DDIM Inversion error can be attributed to the approximation of the diffusion model’s output at step $t\in{1\dots T}$ with the output from step $t-1$ (see eq. 3). Hence, we can define the inversion approximation error for step $t$ as the difference between DM’s output for the target and previous timesteps ${t}$ and ${t-1}$ as:
where $\mathcal{E}_{t}^{S}$ is the true model prediction at step $t$, and $\mathcal{E}_{t}^{I}$ is the inversion’s approximation using the previous step’s output. Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the inversion process. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To ascertain that we measure the error solely for the analyzed step $t$, we start the inversion procedure from the exact latent from step $(t-1)$ (obtained from the sampling path). We split the latent pixels into plain and non-plain areas according to the masks calculated for clean images. More details on the experimental setup can be found in appendix H.
In fig. 4, we present the visualization of calculated differences for each diffusion step. There is a significant difference in the prediction errors for plain and non-plain areas, especially in the initial steps of the noising process. Additionally, in table 6, we show the actual values of the approximation error. By splitting the inversion process into two parts, we can observe that the first 10% of inversion steps contribute more to the overall error than the remaining 90%. Notably, the error in plain areas predominantly accumulates within the first 10% of the inversion steps.
Those observations can be related to recent works [29, 32] that analyze why numerical solvers incur significant errors during the earliest diffusion steps (as $t\rightarrow 0$). Specifically, Lee et al. [29] trace the error to the $1/t$ curvature blow-up of the reverse-time ODE trajectory, whereas Lin et al. [32] attribute the predominance of early DDIM inversion approximations to a singularity arising from commonly used noise schedules. Our experiments extend those studies by showing that the approximation error can mainly be attributed to the plain regions in the original images.
@@ Conclusion.
Early approximations during the DDIM Inversion procedure lead to unequally distributed errors for pixels related to the plain and non-plain image areas, making it the origin of structural patterns and correlations exhibited by the latent encodings.
4 Consequences of the divergence and how to fix them?
After identifying the differences between the Gaussian noises and latents and highlighting the origin of this phenomenon, we finally pose the last question:
Research Question 4: What are the practical consequences of the divergence between noises and inverse DDIM latents, and how to fix them?
Building on top of our findings in section 3.3, which indicated that the initial inversion steps predominantly contribute to the divergence between latent variables and Gaussian noise, we propose to decorrelate the final DDIM latents by replacing the first inversion steps with random Gaussian noise $\epsilon\sim\mathcal{N}(0;\mathcal{I})$ as in the standard forward diffusion process (see appendix I for pseudocode):
$x_{t}=\begin{cases}\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,&\text{if }t\leq t^{\prime}\quad\text{(Forward Diffusion)}\\ \sqrt{\alpha_{t}}{x_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}-\bar{\alpha}_{t}})\cdot\epsilon^{(t)}_{\theta}(x_{t-1},c),&\text{if }t>t^{\prime}\quad\text{(DDIM Inversion)}\end{cases}$ (5)
We first evaluate the effectiveness of this technique, with $N=10000$ images generated with $T=50$ diffusion steps using DiT and IF models. For such generations, we first noise them with a forward diffusion to the intermediate step $t^{\prime}$, followed by the DDIM inversion for $T-t^{\prime}$ steps, starting from the $x_{t^{\prime}}$ and ending up with a new approximation of the starting noise. In table 7, we show that by replacing just $4\%$ of the inversion steps, we can completely remove correlations in the latents, up to the level of random Gaussian noise. Moreover, in table 8, we measure the size of the different parts of images (plain vs non-plain) after saving them with the PNG lossless compression. As is visible, compression is the most effective in the parts related to plain images generated from the DDIM latents, which inclines low diversity of their values. At the same time, replacing only 4% of inversion steps significantly reduces this issue. In the following sections, we showcase several consequences of the divergence between noises and latents and how our simple solution reduces those issues.
Inversion steps DiT IF
replaced by Pixel Reconstruction KL Div. Pixel Reconstruction KL Div.
forward (%$\%T$) Corr. Absolute Error ×$\times 10^{-3}$ Corr. Absolute Error ×$\times 10^{-3}$
Noise ${\mathbf{x_{T}}}$ $0.04$ $0.00$ $0.20$ $0.05$ $0.00$ $0.40$
DDIM latent ${\mathbf{\hat{x}_{T}}}$ $0.16$ $0.05$ $11.57$ $0.64$ $0.07$ $608.25$
$1$ ($2\%$) $0.04$ $0.05$ $0.29$ $0.06$ $0.07$ $0.98$
$1\dots 2$ ($4\%$) $0.04$ $0.07$ $0.25$ $0.05$ $0.07$ $0.48$
$1\dots 5$ ($10\%$) $0.04$ $0.12$ $0.49$ $0.05$ $0.08$ $0.42$
$1\dots 10$ ($20\%$) $0.04$ $0.15$ $0.45$ $0.05$ $0.10$ $0.40$
Table 7: Structures can be removed from DDIM latents by replacing inversion steps with forward diffusion. Using forward diffusion instead of the first $4\%$ of inversion steps brings the resulting latents closer to Gaussian noise without a major degradation in the image reconstruction.
Different prompt generations from:
Model Region Noise Latent Latent
(baseline) DDIM Inv. w/ Forward 4%
IF Plain 17.90 14.92 ( +16.7%) 17.42 ( +2.7%)
Non-plain 18.60 17.35 ( + 6.7%) 18.16 ( +2.4%)
DiT Plain 13.64 11.95 ( +12.4%) 13.27 ( +2.7%)
Non-plain 16.49 15.34 ( + 7.0%) 16.18 ( +1.9%)
Table 8: PNG bit-rate (bits / pixel) after saving only the masked pixels. Higher compression (lower bpp) means less local variability in the pixel stream. Values in parentheses are the percentage change with respect to the noise baseline.
@@ 4.1 Interpolation quality
We start with the task of image interpolation, where the goal, for two given images, is to generate a sequence of semantically meaningful intermediary frames. Numerous methods [8, 50, 47, 63, 60, 1] leverage diffusion models to obtain realistic transitions. Song et al. [50] propose to use the latent variables resulting from the DDIM inversion technique and interpolate them with the spherical linear interpolation (SLERP [48]), that, for two objects $x$ and $y$, with an interpolation coefficient $\lambda\in[0;1]$, is defined as $z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y$, where $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
@@
We compare the quality of interpolations in the noise and latent spaces. To this end, we sample $N=20$k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into the latents. Next, we randomly assign pairs of either two noises or latents, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In fig. 5, we show that, by calculating FID-10k, intermediate noise generations preserve consistent quality along the entire interpolation path – meaning that all interpolated images fall into the real images manifold. In contrast, this property collapses for the setting with latents as inputs. In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall [27], especially in the middle of the interpolation path. Nevertheless, as presented in table 9, replacing only the first step of the DDIM inversion with the forward diffusion significantly reduces the issue, leading almost to the quality of the interpolations between ground-truth noises.
@@
Figure 5: The quality of images generated from spherical linear interpolations of latents ${\mathbf{\hat{x}_{T}}}$ deteriorates along the interpolation path, as indicated by the FID peak in the middle. In contrast, the quality of generations from noise ${\mathbf{x_{T}}}$ interpolations remains stable. Recall metric reveals that the latent interpolation leads to the generation of much less diversity than noise.
4.2 Diversity and quality of image edition
Table 9: Starting inversion with forward diffusion improves quality (FID) of latent interpolations. Replacing first steps improves generations’ diversity (Recall).
Inversion steps DiT
replaced by SLERPλ=0.5
forward (%$\%T$) FID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Noise ${\mathbf{x_{T}}}$ $16.61$ $0.799$ $0.473$
DDIM Latent ${\mathbf{\hat{x}_{T}}}$ $19.68$ $0.795$ $0.415$
$1$ ($2\%$) $16.89$ $0.816$ $0.443$
$1\dots 2$ ($4\%$) $16.74$ $0.817$ $0.443$
$1\dots 5$ ($10\%$) $17.05$ $0.812$ $0.444$
$1\dots 10$ ($20\%$) $17.02$ $0.809$ $0.454$